Solve the following A vertical rectangular surface has a base of 1.2 m and a height of 3 m. If the center of pressure is o.150 m below the center of gravity, how far is the base below the liquid surface 1. h: ? o. 150 m 1-12 t.2 2. A circular gate 1.50 m in diameter is inclined at an angle of 45 Sea water stands on one side of the gate to a height of 10 m above the center of gate. Determine the total force acting perpendicular to the gate and the location. Sea Wa IOm 45 3. A sluice gate is in the form of a circular arc of radius 5 m. Calculate the resultant force acting on the gate as shown below. 5m 4. A sphere 1.5 m in diameter floats half submerged in a tank of Oil (S-0.8) a.) What is the weight of the sphere? b.) What is the minimum weight of an anchor weighing 24 KN/cu.m. that will be required to submerge the sphere completely?
Answer #1

b = 1.2m ak d=3m Yo = h - 1.5 Yp = h - (1.5-0.15) = /h- 1.35) 1.8st is Yp = Ye + In Aye Is = 102x3 = 2.7 m (h-1.35) = (h.1.5)